OpenStudy (anonymous):

Find the slope of the line that contains the points, (1, 2) and (2, 5).

OpenStudy (anonymous):

Part b.) Using the slope that you found in part a and one of the given points, write the equation of the line. Please write your final answer in slope-intercept form.

OpenStudy (saifoo.khan):

slope = y2 - y1 ------ x2 - x1

OpenStudy (anonymous):

To find slope, use the equation x_2-x_1 over y_2-y_1 so we'll come up with 3 for slope

OpenStudy (saifoo.khan):

insert it.

OpenStudy (anonymous):

Err.. other way around for my answer

OpenStudy (anonymous):

Saifoo got creative :D

OpenStudy (saifoo.khan):

YAY! :D

OpenStudy (anonymous):

SO HOW do u do part b??

OpenStudy (anonymous):

Slope intercept form is y=mx+b so insert slope to m and any two points to solve for b

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so i can say y=3x+5??

OpenStudy (anonymous):

We'll use (1,2) and come up with 2=3(1)+b so after solving for b we get b=-1 so take that and plug it into y=mx+b you come up with y=3x-1